OpenStudy (anonymous):

please help Make a conjecture about the next item in the sequence. 6, 2, −4, −8, −14

OpenStudy (anonymous):

6 2 = -4 2 -4 = -6 -4 -8 = -4 -8 -14 = -6 What difference should be next ?

OpenStudy (anonymous):

-18 ??

OpenStudy (anonymous):

Yup , Correct !

OpenStudy (anonymous):

Thanks So much